A rock is dropped from a 300 m –high cliff. How long does it take to fall (a) the first 100 m


and (b) the last 50 m.

1
Expert's answer
2021-11-15T17:54:49-0500

(a) The time is


"t=\sqrt{2h_1/g}=4.5\text{ s}."


(b) The total time of flight is


"T=\sqrt{2H/g}."

The time to cover the first 250 m:


"t_2=\sqrt{2(H-50)/g}."

The time to cover the last 50 m:


"t_3=T-t_2=\sqrt{2H/g}-\sqrt{2(H-50)/g}=0.68\text{ s}."